# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 3.1. Open Kulikov surfaces
Open Kulikov surfaces are normal crossing surfaces whose irreducible components are particular surfaces called 00-surfaces, $1$-surfaces, and $2$-surfaces. We first define below these classes of surfaces.
@@
The fan of the affine toric surface $(\overline{Y},\overline{D})$ is a strictly convex rational cone in $\mathbb{R}^{2}$. A toric 00-surface is then a toric surface whose fan is a refinement of $\sigma$, that is, is obtained by adding rays contained in the interior of $\sigma$. Finally, any 00-surface $(Y,D)$ is obtained from a toric 00-surface $(Y^{\prime},D^{\prime})$ by interior blow-ups. The tropicalization of $(Y,D)$ is obtained by introducing focus-focus singularities moving along the interior rays of the fan of $(Y^{\prime},D^{\prime})$, and pushing them to the origin, as in [40].
@@
A toric $1$-surface is a toric surface whose fan has support a half-plane in $\mathbb{R}^{2}$. Any $1$-surface $(Y,D)$ is obtained from a toric $1$-surface $(Y^{\prime},D^{\prime})$ by interior blow-ups and corner blow-downs. The tropicalization of $(Y,D)$ is obtained by introducing focus-focus singularities moving along the interior rays of the fan of $(Y^{\prime},D^{\prime})$, pushing them until they sit at the origin, as in [40], and eventually removing some interior rays.
@@ 3.1.2. Open Kulikov surfaces
    Let $\{X^{i}\}_{i\in I}$ be the set of irreducible components of $\mathcal{X}_{0}$, and denote by $\{v_{i}\}_{i\in I}$ the corresponding vertices in $B$. For every $i\in I$, denote the union of double curves in $X^{i}$ by $\partial X^{i}=\cup_{j\in I\setminus\{i\}}X^{i}\cap X^{j}$. Then, for each interior vertex, $v_{i}\in\mathrm{Int}(B)$, the pair $(X^{i},\partial X^{i})$ is a 2-surface, and for each boundary vertex, $v_{i}\in\partial B$, the pair $(X^{i},\partial X^{i})$ is either a 00-surface or a $1$-surface.
@@
If one replaces i) in Definition 3.7 by the condition that $B$ is a sphere, and ii) by the condition that $(X^{i},\partial X^{i})$ is a $2$-surface for every vertex $v_{i}$ of $B$, then one obtains the definition of a Type III Kulikov surface, appearing in the context of Type III Kulikov degenerations of K3 surfaces [7, 35, 69, 80]. On the other hand, if one still replaces i) by the condition that $B$ is a sphere, but if one replaces ii) by the condition that every $(X^{i},\partial X^{i})$ is a $2$-surface, except one which is a Hirzebruch-Inoue surface, then one obtains the definition of a “Type III anticanonical pair”, appearing in the context of smoothings of cusp singularities in [30, 31].
@@ 3.2. Affinization of open Kulikov surfaces
$\overline{\mathcal{X}}_{0}:=\bigcup_{v\in P_{\mathbb{Z}}(0)}\overline{X}^{v}=\bigcup_{i\in\mathbb{Z}/N\mathbb{Z}}\overline{X}^{v_{i}}$ (3.1)
by gluing as follows the affine toric surfaces $\overline{X}^{v_{i}}$. For every $i\in\mathbb{Z}/N\mathbb{Z}$, we identify $\overline{X}^{v_{i}}$ and $\overline{X}^{v_{i+1}}$ along their toric divisors $(\partial\overline{X}^{v_{i}})_{i+1}\simeq\mathbb{A}^{1}$ and $(\partial\overline{X}^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$ by the isomorphism induced by the gluing isomorphism between the irreducible components $(\partial{X}^{v_{i}})_{i+1}$ and $(\partial{X}^{v_{i+1}})_{i}$ of the double locus $\mathcal{X}_{0}$ corresponding to the two boundary edges of $\mathscr{P}$ adjacent to $v_{i}$. In particular, this isomorphism preserves $0\in\mathbb{A}^{1}$, and so identifies the torus fixed points of $\overline{X}^{v_{i}}$ and $\overline{X}^{v_{i+1}}$.
By construction, the irreducible components of $\overline{\mathcal{X}}_{0}$ are the the affine toric surfaces $\overline{X}^{v_{i}}$, and so $\overline{\mathcal{X}}_{0}$ is itself affine by [50, Exercise III.3.2]. Moreover, the double locus of $\overline{\mathcal{X}}_{0}$ is a union of copies of affine lines $(\partial X^{v_{i}})_{i+1}\simeq(\partial X^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$, and all intersecting at a single common point $x_{0}\in\overline{\mathcal{X}}_{0}$, referred below as the central point of $\overline{\mathcal{X}}_{0}$, see Figure 3.1. It follows that $\overline{\mathcal{X}}_{0}$ is a degenerate cusp singularity as in [85, §1], and so is in particular Gorenstein semi-log-canonical by [65, Theorem 4.21].
@@
Finally, we construct a morphism
$f_{0}:\mathcal{X}_{0}\longrightarrow\overline{\mathcal{X}}_{0}\,,$ (3.2)
as follows. We first define a morphism $f^{v}:X^{v}\rightarrow\overline{\mathcal{X}}_{0}$ for every integral point $v\in P_{\mathbb{Z}}$. To do that, we distinguish the three cases of 00-, $1$-, and $2$-surface:
@@
For every $v\in(\partial P)_{\mathbb{Z}}$, the affinization map $X^{v}\rightarrow\overline{X}^{v}$ is proper and surjective, and so it follows from the definition of $f_{0}$ that $f_{0}$ is also proper and surjective. To compute $R^{i}(f_{0})_{\star}\mathcal{O}_{\mathcal{X}_{0}}$, one uses the resolution of $\mathcal{O}_{\mathcal{X}_{0}}$ given by the complex
@@
We have
$\overline{\mathcal{X}}_{0}=\mathrm{Spec}\,H^{0}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})\,,$
and $f_{0}:\mathcal{X}_{0}\longrightarrow\overline{\mathcal{X}}_{0}$ is the affinization morphism of $\mathcal{X}_{0}$.
@@ 3.3. $d$-semistable and generic open Kulikov surfaces
Recall from [34] that a normal crossing surface $X$, with singular locus $X_{\mathrm{sing}}$, is called $d$-semistable if $\mathcal{E}xt^{1}(\Omega_{X},\mathcal{O}_{X})\cong{\mathcal{O}_{X}}_{\mathrm{sing}}$. To study $d$-semistability of open Kulikov surfaces, we adapt to open Kulikov surfaces the study of Type III Kulikov surfaces done in [7, §7.B]
@@
Define $\xi_{i}:=\sum_{j}(D_{ij}-D_{ji})\in\widetilde{\mathbb{L}}$. For a $\psi\in\mathrm{Hom}(\widetilde{\mathbb{L}},\mathbb{C}^{\star})$, we define
$\psi_{i}:=\psi(\xi_{i})\in\mathbb{C}^{\star}\,.$
Note that when $\mathcal{X}_{0}$ admits a smoothing $\mathcal{X}$, then $\xi_{i}$ is the class of the line bundle $\mathcal{O}_{\mathcal{X}}(-V_{i})|_{\mathcal{X}_{0}}$ in $\mathrm{Pic}(\mathcal{X}_{0})=\widetilde{\mathbb{L}}$.
@@
The normal crossing surface $\mathcal{X}_{0,\psi}$ is $d$-semistable if $T^{1}_{\mathcal{X}_{0,\psi}}=\mathcal{E}xt^{1}(\Omega^{1}_{\mathcal{X}_{0,\psi}},\mathcal{O}_{\mathcal{X}_{0,\psi}})\simeq\mathcal{O}_{D}$, where $D$ is the double locus of $\mathcal{X}_{0,\psi}$, see [33, 35]. The sheaf $T^{1}_{\mathcal{X}_{0,\psi}}$ is an invertible sheaf on $D$ and its class lies in $\operatorname{Pic}^{0}(D)=\mathrm{Hom}(H_{1}(\Gamma(D),\mathbb{C}^{*}))\simeq(\mathbb{C}^{*})^{g}$, where $\Gamma(D)$ is the dual graph of $D$ and $g$ is its genus. The group $H_{1}(\Gamma(D)\simeq\mathbb{Z}^{g}$ as a basis has loops $\Gamma(D_{i})$ for the boundaries of the compact components $V_{i}$. The restriction $T^{1}_{\mathcal{X}_{0,\psi}}|_{D_{i}}$ is $\psi_{i}\in\operatorname{Pic}^{0}(D_{i})=\mathbb{C}^{*}$. It follows that $T^{1}_{\mathcal{X}_{0,\psi}}=1$ in $\operatorname{Pic}^{0}(D)$ if and only if all $\psi_{i}=1$. ∎
It follows from Lemma 3.13 that $\mathcal{X}_{0,\psi}$ is $d$-semistable if and only if $\psi$ is contained in the torus $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})\subset\mathrm{Hom}(\widetilde{\mathbb{L}},\mathbb{C}^{\star})$, where $\mathbb{L}$ is the the quotient of $\widetilde{\mathbb{L}}$ by the span of the classes $\xi_{i}$ such that $V_{i}$ is compact.
We now study the effect of a change of gluing of the irreducible components on $\psi\in\mathrm{Hom}(\widetilde{\mathbb{L}},\mathbb{C}^{\star})$.
@@
For every $\mathcal{L}\in\widetilde{\mathbb{L}}$, changing the gluing along $D_{ij}=V_{i}\cap V_{j}$ by $c_{ij}\in\mathbb{C}^{\star}$ changes $\psi(\mathcal{L})$ by
@@
This follows from [7, Theorem 7.9] by tracing the quasi-isomorphism between the Picard complex of and the gluing complex, which gives an explicit embedding of $\mathbb{L}$ into $\mathbb{Z}^{E}\oplus\mathbb{Z}^{Q}$, where $E=\{D_{ij}\mid i<j\}$ is the set of edges in the dual complex of $\mathcal{X}_{0}$, and $Q$ is the set of internal blowups. Changing the gluings acts on $\mathrm{Hom}(\mathbb{Z}^{E},\mathbb{C}^{*})$ in an obvious way, as $\prod_{i<j}c_{ij}^{\mathcal{L}\cdot D_{ij}}$. ∎
@@
Consider a component $D_{ij}=V_{i}\cap V_{j}$ of the double locus of $\mathcal{X}_{0}$ which is not nodal genus one curve. Let $V_{k}$ and $V_{l}$ the two irreducible components of $\mathcal{X}_{0}$ such that $D_{ijk}:=D_{ij}\cap V_{k}$ and $D_{ijl}:=D_{ij}\cap V_{l}$ are the two triple points of $\mathcal{X}_{0}$ on $D_{ij}$. Then, changing the gluing along $D_{ij}$ by $c_{ij}\in\mathbb{C}^{\star}$ changes the periods $\psi_{m}$ as follows:
@@
By [34, Lemma 1.6], arbitrary moving the position of the internal blow-ups of the fixed toric model produces a universal family $(V_{i}^{\mathrm{univ}},V_{i}^{\mathrm{univ}})\rightarrow\mathrm{Hom}(\mathbb{L}_{i},\mathbb{C}^{\star})$, where every log Calabi–Yau pair deformation equivalent to $(Y,D)$ appears. There is a natural map $\mathbb{L}_{i}\rightarrow\widetilde{\mathbb{L}}$, and so, by composition with the quotient map $\widetilde{\mathbb{L}}\rightarrow\mathbb{L}$, a map $\mathbb{L}_{i}\rightarrow\mathbb{L}$, and so a map of tori $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})\rightarrow\mathrm{Hom}(\mathbb{L}_{i},\mathbb{C}^{\star})$, sending a $d$-semistable deformation of $\mathcal{X}_{0}$ on the corresponding deformation of the irreducible component $(V_{i},D_{i})$.
@@
For every irreducible component $(V_{i},D_{i})$ of $\mathcal{X}_{0}$, defined $d(i)$ as the smallest nonnegative integer such that there exists a sequence of irreducible components $V_{j_{0}}$, $V_{j_{1}}$,…, $V_{j_{d(i)}}$ such that $\dim V_{j_{k}}\cap V_{j_{k+1}}=1$ for all $k$, $V_{j_{0}}=V_{i}$, and $V_{j_{d(i)}}$ is non-compact. For example, we have $d(i)=0$ if and only if $V_{i}$ is non-compact. In other words, $d(i)$ is the distance following edges in the dual intersection complex $(P,\mathscr{P})$ of $\mathcal{X}_{0}$ between the vertex corresponding to $v_{i}$ and the boundary of $P$.
Start with any $\mathcal{X}_{0}=\cup_{i}(V_{i},D_{i})$. We will show that one can arbitrary deform the components $(V_{i},D_{i})$ while rescaling the gluings along the double curves to make $\mathcal{X}_{0}$ $d$-semistable. More precisely, we prove by descending induction on $d$ that one can arbitrary deform the components $(V_{i},D_{i})$ with $d(i)\geq d$ while rescaling the gluings along the double curves to keep $\psi_{i}=1$ for all $i$ such that $V_{i}$ is compact and $d(i)\geq d$. This implies Theorem 3.16 when $d=0$.
Since $\mathcal{X}_{0}$ has finitely many irreducible components, the set of values of $d(i)$ is finite, and so the result is trivially true for $d$ large enough. For the induction step, assume that the result holds for $d+1$, and let us show that it holds for $d$.
Let $V_{j}$ be a component with $d(j)=d-1$. If $d-1=0$, let $D_{jk_{1}},\cdots,D_{jk_{m}}$ be the chain of irreducible components of $D_{j}$, such that $D_{jk_{l}}\cap D_{jk_{l+1}}\neq\emptyset$. If $d-1>0$, then there is at least one component $D_{jk_{0}}$ of $D_{j}$ such that $d(k_{0})=d-2$, and denote by $D_{jk_{1}},\cdots,D_{jk_{m}}$ be the chain of the remaining irreducible components of $D_{j}$ – see Figure 3.2. Note that, for every $1\leq l\leq m$, we have $d(k_{l})\leq d$.
@@
We apply the previous process to every component $V_{j}$ with $d(j)=d-1$. Every component $V_{i}$ with $d(i)=d$ is adjacent to a component $V_{j}$ with $d(j)=d-1$, so appears in that way, and is ultimately deformed while $\psi_{i}$ is set equal to $1$. In this process, we only change the gluing along double curves of the form $D_{ij}$, with triple points $D_{ijk}$ and $D_{ijl}$, such that $d(i)$, $d(j)$, $d(k)$, or $d(l)$ are all equal to $d-2$, $d-1$, or $d$. In particular, by Lemma 3.15, the values of $\psi_{i}$ with $d(i)>d$ are never changed, and so stay equal to their value $1$ given by the induction hypothesis. ∎
As in [41, Definition 1.4], we call a (possibly non-compact) log Calabi–Yau surface generic if it does not contain any internal $(-2)$-curve. We define below a notion of “generic” for open Kulikov surfaces. To motivate the definition, we first introduce the notions of M1 and M2 modifications of an open Kulikov surface, following the classical case of Type III Kulikov surfaces – e.g. see [7, Definition 7.15]. If $\mathcal{X}_{0}$ is an open Kulikov degeneration, and $C$ an internal exceptional curve contained in a component $V_{i}$, intersecting the double locus $V_{i}\cap V_{j}$ at the point $p$, then, blowing-down $C$ in $V_{i}$, and blowing up $p$ in $V_{j}$ produces a new open Kulikov surface $\mathcal{X}_{0}^{\prime}$, called the M1 modification of $\mathcal{X}_{0}$ along $C$. Similarly, if $C=V_{i}\cap V_{j}$ is an exceptional component of the double locus of $\mathcal{X}_{0}$, containing the triple points $p_{ijk}=V_{i}\cap V_{j}\cap V_{k}$ and $p_{ijl}=V_{i}\cap V_{j}\cap V_{l}$, then, blowing-down $C$ in $V_{i}$ and $V_{j}$, and blowing-up $p_{ijk}$ and $p_{ijl}$ in $V_{k}$ and $V_{l}$ respectively, produces a new open Kulikov surface $\mathcal{X}_{0}^{\prime}$, called the M2 modification of $\mathcal{X}_{0}$ along $C$. Under an M1 modification, an internal $(-2)$-curve can disappear and be replaced by a particular configuration of curves – see Figure 3.3, that we call degenerate $(-2)$-curves and define as follows.

### 30 Mar 2024 v2

@@ 3.1. Open Kulikov surfaces
Open Kulikov surfaces are normal crossing surfaces whose irreducible components are particular types of surfaces called 00-surfaces, $1$-surfaces, and $2$-surfaces, defined as below.
@@
The fan of the affine toric surface $(\overline{Y},\overline{D})$ is a strictly convex rational cone $\sigma$ in $\mathbb{R}^{2}$. A toric 00-surface is then a toric surface whose fan is a refinement of $\sigma$. Finally, any 00-surface $(Y,D)$ is obtained from a toric 00-surface $(Y^{\prime},D^{\prime})$ by interior blow-ups. The tropicalization of $(Y,D)$ is obtained by introducing focus-focus singularities moving along the interior rays of the fan of $(Y^{\prime},D^{\prime})$, and pushing them to the origin, as in [41].
@@
A toric $1$-surface is a toric surface whose fan has support a half-plane in $\mathbb{R}^{2}$. Any $1$-surface $(Y,D)$ is obtained from a toric $1$-surface $(Y^{\prime},D^{\prime})$ by interior blow-ups and corner blow-downs. The tropicalization of $(Y,D)$ is obtained by introducing focus-focus singularities moving along the interior rays of the fan of $(Y^{\prime},D^{\prime})$, pushing them to the origin, as in [41], and eventually removing some interior rays.
@@ 3.1.2. Open Kulikov surfaces
    Let $\{X^{i}\}_{i\in I}$ be the set of irreducible components of $\mathcal{X}_{0}$, and denote by $\{v_{i}\}_{i\in I}$ the corresponding vertices in $B$. For every $i\in I$, denote the union of double curves in $X^{i}$ by $\partial X^{i}=\cup_{j\in I\setminus\{i\}}X^{i}\cap X^{j}$. Denote by $\tilde{X}^{i}$ the normalization of $X^{i}$ and by $\partial\tilde{X}^{i}$ the pre-image of $\partial X^{i}$ in $\tilde{X}^{i}$. Then, for each interior vertex $v_{i}\in\mathrm{Int}(B)$, the pair $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is a 2-surface, and for each boundary vertex $v_{i}\in\partial B$, the pair $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is either a 00-surface or a $1$-surface.
@@
If one replaces i) in Definition 3.7 by the condition that $B$ is a sphere, and ii) by the condition that $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is a $2$-surface for every vertex $v_{i}$ of $B$, then one obtains the definition of a Type III Kulikov surface, appearing in the context of Type III Kulikov degenerations of K3 surfaces [7, 35, 70, 81]. On the other hand, if one still replaces i) by the condition that $B$ is a sphere, but one rather replaces ii) by the condition that every $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is a $2$-surface, except one which is a Hirzebruch-Inoue surface, then one obtains the definition of a “Type III anticanonical pair”, appearing in the context of smoothings of cusp singularities in [30, 31].
@@ 3.2. Affinization of open Kulikov surfaces
$\overline{\mathcal{X}}_{0}:=\bigcup_{v\in P_{\mathbb{Z}}(0)}\overline{X}^{v}=\bigcup_{i\in\mathbb{Z}/N\mathbb{Z}}\overline{X}^{v_{i}}\,,$ (3.1)
where for every $i\in\mathbb{Z}/N\mathbb{Z}$, we identify $\overline{X}^{v_{i}}$ and $\overline{X}^{v_{i+1}}$ along their toric divisors $(\partial\overline{X}^{v_{i}})_{i+1}\simeq\mathbb{A}^{1}$ and $(\partial\overline{X}^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$ by the isomorphism induced by the gluing isomorphism between the irreducible components $(\partial{X}^{v_{i}})_{i+1}$ and $(\partial{X}^{v_{i+1}})_{i}$ of the double locus $\mathcal{X}_{0}$ corresponding to the two boundary edges of $\mathscr{P}$ adjacent to $v_{i}$. In particular, this isomorphism preserves $0\in\mathbb{A}^{1}$, and so identifies the torus fixed points of $\overline{X}^{v_{i}}$ and $\overline{X}^{v_{i+1}}$.
By construction, the irreducible components of $\overline{\mathcal{X}}_{0}$ are the the affine toric surfaces $\overline{X}^{v_{i}}$, and so $\overline{\mathcal{X}}_{0}$ is itself affine by [51, Exercise III.3.2]. Moreover, the double locus of $\overline{\mathcal{X}}_{0}$ is a union of copies of affine lines $(\partial X^{v_{i}})_{i+1}\simeq(\partial X^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$, and all intersecting at a single common point $x_{0}\in\overline{\mathcal{X}}_{0}$, referred below as the central point of $\overline{\mathcal{X}}_{0}$, see Figure 3.1. It follows that $\overline{\mathcal{X}}_{0}$ is a degenerate cusp singularity as in [86, §1], and so is in particular Gorenstein semi-log-canonical by [68, Theorem 4.21].
@@
Finally, we construct a morphism $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ as follows. We first define a morphism $f^{v}:X^{v}\rightarrow\overline{\mathcal{X}}_{0}$ for every integral point $v\in P_{\mathbb{Z}}$. To do that, we distinguish the three cases of 00-, $1$-, and $2$-surface:
@@
For every $v\in(\partial P)_{\mathbb{Z}}$, the affinization map $X^{v}\rightarrow\overline{X}^{v}$ is proper and surjective, and so $f_{0}$ is also proper and surjective. To compute $R^{i}(f_{0})_{\star}\mathcal{O}_{\mathcal{X}_{0}}$, one uses the resolution of $\mathcal{O}_{\mathcal{X}_{0}}$ given by the complex
@@
We have $\overline{\mathcal{X}}_{0}=\mathrm{Spec}\,H^{0}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})$, and $f_{0}:\mathcal{X}_{0}\longrightarrow\overline{\mathcal{X}}_{0}$ is the affinization morphism of $\mathcal{X}_{0}$.
@@ 3.3. $d$-semistable and generic open Kulikov surfaces
Recall from [34] that a normal crossing surface $X$, with singular locus $X_{\mathrm{sing}}$, is called $d$-semistable if $\mathcal{E}xt^{1}(\Omega_{X},\mathcal{O}_{X})\cong{\mathcal{O}_{X}}_{\mathrm{sing}}$. We will study $d$-semistability of open Kulikov surfaces, similarly as in [7, §7.B], where $d$-semistability of Type III Kulikov surfaces is studied.
@@
Define $\xi_{i}:=\sum_{j}(D_{ij}-D_{ji})\in\widetilde{\mathbb{L}}$. For example, when $\mathcal{X}_{0}$ admits a smoothing $\mathcal{X}$, $\xi_{i}$ is the class of the line bundle $\mathcal{O}_{\mathcal{X}}(-V_{i})|_{\mathcal{X}_{0}}$ in $\mathrm{Pic}(\mathcal{X}_{0})=\widetilde{\mathbb{L}}$. In general, for every element $\psi\in\mathrm{Hom}(\widetilde{\mathbb{L}},\mathbb{C}^{\star})$, we define $\psi_{i}:=\psi(\xi_{i})\in\mathbb{C}^{\star}$.
@@
The normal crossing surface $\mathcal{X}_{0,\psi}$ is $d$-semistable if $T^{1}_{\mathcal{X}_{0,\psi}}=\mathcal{E}xt^{1}(\Omega^{1}_{\mathcal{X}_{0,\psi}},\mathcal{O}_{\mathcal{X}_{0,\psi}})\simeq\mathcal{O}_{D}$, where $D$ is the double locus of $\mathcal{X}_{0,\psi}$, see [33, 35]. The sheaf $T^{1}_{\mathcal{X}_{0,\psi}}$ is an invertible sheaf on $D$ and its class lies in $\operatorname{Pic}^{0}(D)=\mathrm{Hom}(H_{1}(\Gamma(D),\mathbb{C}^{*}))\simeq(\mathbb{C}^{*})^{g}$, where $\Gamma(D)$ is the dual graph of $D$ and $g$ is its genus. The group $H_{1}(\Gamma(D))$ is isomorphic to $\mathbb{Z}^{g}$, with a basis consisting of the loops $\Gamma(D_{i})$ corresponding to the boundaries of the compact components $V_{i}$. The restriction $T^{1}_{\mathcal{X}_{0,\psi}}|_{D_{i}}$ is $\psi_{i}\in\operatorname{Pic}^{0}(D_{i})=\mathbb{C}^{*}$. It follows that $T^{1}_{\mathcal{X}_{0,\psi}}=1$ in $\operatorname{Pic}^{0}(D)$ if and only if all $\psi_{i}=1$. ∎
It follows from Lemma 3.13 that $\mathcal{X}_{0,\psi}$ is $d$-semistable if and only if $\psi$ is contained in the torus $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})\subset\mathrm{Hom}(\widetilde{\mathbb{L}},\mathbb{C}^{\star})$, where $\mathbb{L}$ is the the quotient of $\widetilde{\mathbb{L}}$ by the span of the classes $\xi_{i}$ such that $V_{i}$ is compact. The following result shows how $\psi$ changes by changing the gluing of the irreducible components.
@@
For every $\mathcal{L}\in\widetilde{\mathbb{L}}$, changing the gluing along $D_{ij}=V_{i}\cap V_{j}$ by multiplication by $c_{ij}\in\mathbb{C}^{\star}$ changes $\psi(\mathcal{L})$ by
@@
This follows from [7, Theorem 7.9] by tracing the quasi-isomorphism between the Picard complex of and the gluing complex, which gives an explicit embedding of $\mathbb{L}$ into $\mathbb{Z}^{E}\oplus\mathbb{Z}^{Q}$, where $E=\{D_{ij}\mid i<j\}$ is the set of edges in the dual complex of $\mathcal{X}_{0}$, and $Q$ is the set of internal blowups. Changing the gluings acts on $\mathrm{Hom}(\mathbb{Z}^{E},\mathbb{C}^{*})$ in an obvious way, by multiplication by $\prod_{i<j}c_{ij}^{\mathcal{L}\cdot D_{ij}}$. ∎
@@
Consider a component $D_{ij}=V_{i}\cap V_{j}$ of the double locus of $\mathcal{X}_{0}$ which is not nodal genus one curve. Let $V_{k}$ and $V_{l}$ the two irreducible components of $\mathcal{X}_{0}$ such that $D_{ijk}:=D_{ij}\cap V_{k}$ and $D_{ijl}:=D_{ij}\cap V_{l}$ are the two triple points of $\mathcal{X}_{0}$ on $D_{ij}$. Then, changing the gluing along $D_{ij}$ by multiplication by $c_{ij}\in\mathbb{C}^{\star}$ changes the periods $\psi_{m}$ as follows:
@@
By [34, Lemma 1.6], arbitrary moving the position of the internal blow-ups of the fixed toric model produces a universal family $(V_{i}^{\mathrm{univ}},V_{i}^{\mathrm{univ}})\rightarrow\mathrm{Hom}(\mathbb{L}_{i},\mathbb{C}^{\star})$, where every log Calabi–Yau pair deformation equivalent to $(Y,D)$ appears. The map $\mathbb{L}_{i}\rightarrow\mathbb{L}$, obtained by composition of the natural map $\mathbb{L}_{i}\rightarrow\widetilde{\mathbb{L}}$ with the quotient map $\widetilde{\mathbb{L}}\rightarrow\mathbb{L}$, induces a map of tori $\mathrm{Hom}(\mathbb{L},\mathbb{C}^{\star})\rightarrow\mathrm{Hom}(\mathbb{L}_{i},\mathbb{C}^{\star})$, sending a $d$-semistable deformation of $\mathcal{X}_{0}$ on the corresponding deformation of the irreducible component $(V_{i},D_{i})$.
@@
For every irreducible component $(V_{i},D_{i})$ of $\mathcal{X}_{0}$, define $d(i)$ as the smallest nonnegative integer such that there exists a sequence of irreducible components $V_{j_{0}}$, $V_{j_{1}}$,…, $V_{j_{d(i)}}$ such that $\dim V_{j_{k}}\cap V_{j_{k+1}}=1$ for all $k$, $V_{j_{0}}=V_{i}$, and $V_{j_{d(i)}}$ is non-compact. For example, we have $d(i)=0$ if and only if $V_{i}$ is non-compact. In other words, $d(i)$ is the distance following edges in the dual intersection complex $(P,\mathscr{P})$ of $\mathcal{X}_{0}$ between the vertex corresponding to $v_{i}$ and the boundary of $P$.
Start with any $\mathcal{X}_{0}=\cup_{i}(V_{i},D_{i})$. We show below that one can arbitrary deform the components $(V_{i},D_{i})$ while rescaling the gluings along the double curves to make $\mathcal{X}_{0}$ $d$-semistable. More precisely, we prove by descending induction on $d$ that one can arbitrary deform the components $(V_{i},D_{i})$ with $d(i)\geq d$ while rescaling the gluings along the double curves to keep $\psi_{i}=1$ for all $i$ such that $V_{i}$ is compact and $d(i)\geq d$. This implies Theorem 3.16 when $d=0$.
Since $\mathcal{X}_{0}$ has finitely many irreducible components, the set of values of $d(i)$ is finite, and so the result is immediate for $d$ large enough. For the induction step, assume that the result holds for $d+1$. In what follows, we show that it holds for $d$.
Let $V_{j}$ be a component with $d(j)=d-1$. If $d-1=0$, let $D_{jk_{1}},\cdots,D_{jk_{m}}$ be the chain of irreducible components of $D_{j}$, such that $D_{jk_{l}}\cap D_{jk_{l+1}}\neq\emptyset$. If $d-1>0$, then there is at least one component $D_{jk_{0}}$ of $D_{j}$ such that $d(k_{0})=d-2$, and denote by $D_{jk_{1}},\cdots,D_{jk_{m}}$ the chain of the remaining irreducible components of $D_{j}$ – see Figure 3.2. Note that, for every $1\leq l\leq m$, we have $d(k_{l})\leq d$.
@@
We apply the previous process to every component $V_{j}$ with $d(j)=d-1$. Every component $V_{i}$ with $d(i)=d$ is adjacent to a component $V_{j}$ with $d(j)=d-1$, and so is ultimately deformed while $\psi_{i}$ is set equal to $1$. In this process, we only change the gluing along double curves of the form $D_{ij}$, with triple points $D_{ijk}$ and $D_{ijl}$, such that $d(i)$, $d(j)$, $d(k)$, or $d(l)$ are all equal to $d-2$, $d-1$, or $d$. In particular, by Lemma 3.15, the values of $\psi_{i}$ with $d(i)>d$ are never changed, and so stay equal to their value $1$ given by the induction hypothesis. ∎
As in [42, Definition 1.4], we call a (possibly non-compact) log Calabi–Yau surface generic if it does not contain any internal $(-2)$-curve. In what follows, we define generic open Kulikov surfaces. To motivate the definition, we first introduce the notions of M1 and M2 modifications of an open Kulikov surface, following the classical case of Type III Kulikov surfaces – e.g. see [7, Definition 7.15]. If $\mathcal{X}_{0}$ is an open Kulikov surface, and $C$ an internal exceptional curve contained in a component $V_{i}$, intersecting the double locus $V_{i}\cap V_{j}$ at the point $p$, then, blowing-down $C$ in $V_{i}$, and blowing up $p$ in $V_{j}$ produces a new open Kulikov surface, called the M1 modification of $\mathcal{X}_{0}$ along $C$. Similarly, if $C=V_{i}\cap V_{j}$ is an exceptional component of the double locus of $\mathcal{X}_{0}$, containing the triple points $p_{ijk}=V_{i}\cap V_{j}\cap V_{k}$ and $p_{ijl}=V_{i}\cap V_{j}\cap V_{l}$, then, blowing-down $C$ in $V_{i}$ and $V_{j}$, and blowing-up $p_{ijk}$ and $p_{ijl}$ in $V_{k}$ and $V_{l}$ respectively, produces a new open Kulikov surface, called the M2 modification of $\mathcal{X}_{0}$ along $C$. Under an M1 modification, an internal $(-2)$-curve can disappear and be replaced by a particular configuration of curves – see Figure 3.3, that we call degenerate $(-2)$-curves and define as follows.